# Toward Human-Quantum Computer Interaction: Interface Techniques for Usable Quantum Computing

### v2

@@ 3.1. Interfaces for Quantum Computing
Prior work in understanding how people use QC tools sets useful context for our work. By interviewing seven scientists in different domains, Ashktorab et al. (ashktorab2019:hqci) report the needs for assisting beginners with using QC tools with limited mathematical knowledge. They also note that after passing a certain point in training, one might need and want to move on to more professional tools for a higher degree of freedom. Furthermore, they recognize the importance of sharing and replicating quantum programs as well as visualizing quantum information in a scalable way given that $n$ qubits result in $2^{n}$ states (worst case). Prior work based on education settings (mykhailova2022:softwareQC; hu2024:qcLearning; tappert2019:experience; salehi2022:csQC) focused on the implementation of simpler circuits representing basic concepts, such as superposition and entanglement, and provided conceptual understanding of complex algorithms like quantum chemistry or optimization problems. In addition, by analyzing self-paced interactive textbook, Wootton et al. (wootton2021:interactiveTextbook) found that learners tended to stop at earlier chapters and frequently referred to appendix for linear algebra. These findings indicate two competing goals for QC support: introducing complex quantum algorithms in a conceptual way vs. training learners to work with low-level components of those algorithms. This work proposes and demonstrates approaches to conceptually representing components of quantum algorithms while not blocking the degree of freedom offered by low-level toolkits.
@@
Visualization takes an important role in making QC more understandable (bethel2023:qcvis), such as Bloch sphere (bloch1946) for qubit states, circuit diagrams, and measurement histograms for QC output. On top of those common visual representations, prior approaches focused on interactive interfaces for circuit composition and result analysis. First, interactive circuit composers like IBM Quantum Composer (ibmComposer) and Quirk (quirk) let users drag and drop gates on a circuit visualization. Yet, those tools still require a deeper understanding of how each gate functions with respect to bigger conceptual ideas, which Classiq (classiq) is trying to address with an interactive whiteboard (Figure 2A). In addition, to improve the interpretability of changes to qubits’ states along with operations in a circuit, prior work proposed approaches to integrating this information on a circuit diagram. For instance, QuFlow (lin2018:quflow) shows possible states of qubits with their probabilities in number on a circuit, and QuantumEyes (ruan2024:quantumeyes) shows them as an area chart with small diagrams explaining how each gate affects the probabilities of qubit states. VENUS (ruan2023:venus) adopts geometrical probability representations for multi-qubit states given that traditional 3D representations like Bloch spheres are limited to show multi-qubit states effectively. However, showing qubit state probabilities on a circuit are feasible only when a circuit includes a few qubits because probability information is not tractable for circuits with more than a few qubits, which is why we need QC. Instead, Quantivine (wen2024:quantivine) abstracts a circuit representation in terms of structural patterns (e.g., repeated gates) and allows for filtering a few qubits for more details like their interactions with other qubits. Note that one can easily encode the structure of a circuit using Qiskit and check their decompositions. Taking one step further, our work proposes and implements interactive visualizations for comparing logical and physical circuits and inspecting how physical qubits’ status change over the execution of a circuit.
@@
In analyzing QC results, developers need to consider non-negligible amount of gate and readout errors as well as adapt the histogram-based outputs for specific problem types. To enable that, VACSEN (ruan2023:vacsen) parallelly shows quantum circuits with the decoherence time and error rates of the gates and measurements involved in executing the circuits. In particular, VACSEN shows error information with timestamps because QC machines’ physical properties change over time. To better support QC-based machine learning (quantum machine learning or QML), VIOLET (ruan2024:violet) offers a expert-targeted dashboard that shows QC outputs in the context of neural network (e.g., output states as features that a model has learned). Yet, prior work does not offer a direct way to represent QC outputs with error rates and/or in a human-understandable format, to which our work aims to propose approaches to complement.
@@
In adopting these complex computing approaches, developers frequently use computational notebooks (e.g., Jupyter Notebook) as they are easy to share, reaccess (e.g., compared to running them on a terminal), and refine (wang2024:supernova). Computational notebooks allow for integrating custom widgets, making it easier to combine work pipelines (wang2024:supernova; harden2024:hni). Yet, prior approaches in QC and complex computing methods tend to showcase standalone tools than those integrated with working environments like computational notebooks. In addition, they often lack support for tasks pertaining to scripting and sharing codes for using them. To address these integration-wise challenges, our work aims to propose interaction techniques for a wider range of QC lifecycle by integrating them within computational notebooks.
@@ 3.2. Interface Support for Data-intensive Work
Prior work has proposed interfaces for different tasks in data-intensive work pipelines, such as data analysis and statistical modeling. Given data work is often done on a Jupyter notebook environment, for example, prior work suggested widgets for various tasks like developing statistical models (jun2022:tisane) and analyzing text data (lam2024:lloom). Dashboard-based interfaces are also common for complex statistical analysis (liu2021:boba; sarma2024:milliways) and literature analysis (narechania2022:vitality). Those interfaces commonly offer descriptions regarding steps that a user is taking. On the other hand, current QC users need to use the same tool regardless of their tasks because many tasks are constrained by QPUs provided via those tools. For instance, gates defined using Cirq can be optimized only by Cirq because the optimization algorithms rely on QPUs accessed via Cirq. Furthermore, users need to frequently switch between the above described QC tasks, requiring interfaces overarching those tasks. Thus, our work also explores how to adapt prior user interface techniques across various QC tasks.
@@ 4.2. Identified Challenges
(C1) Difficulty in expressing ideas and interpreting results conceptually. A recurring topic during our design iterations was the unmatched connection between conceptual ideas and quantum information. In both theoretical and applied QC courses, for example, Kim learned Shor’s algorithm (shors:algorithm) for factoring a large natural number into two other numbers, which is expected to advance quantum cryptography. However, he learned only about theoretical proofs and potential application areas (which is common in many QC classes (hu2024:qcLearning; tappert2019:experience)), but was not able to apply the algorithm to novel problems. It is tricky to try out different numbers for Shor’s algorithm because doing so requires more profound knowledge than what typical beginners would have. Qiskit provides a tutorial for how to do so [^5] https://github.com/qiskit-community/qiskit-community-tutorials/blob/master/algorithms/shor_algorithm.ipynb, but it only shows a case where one factor is fixed at 15. To apply a different factor, Kim had to come up with a method to encode that number using lots of X and Swap gates [^6] A Swap gate changes the states of two qubits, which was beyond Kim’s initial capability (c.f. (beauregard2003:shors)).
Similarly, when Jeng was developing image-convolution techniques for a quantum machine learning (QML) project, she had to come up with a relatively large suite of custom codes to convert images to quantum states (and vice versa). Currently, QML developers lack shared standards that work generically with other image-processing algorithms. Instead, they often need to edit codes from tutorials, making it tedious to apply those codes for novel use cases. Even though more experienced in QC research, Smith was not necessarily able to immediately come up with those methods because Jeng focused on specific QML techniques while Smith was more interested scalable optimization techniques across different application scenarios.
@@ 5. Design Principles for QC Interfaces
Based on our iterative design process and prior work in QC interfaces (ashktorab2019:hqci; lin2018:quflow; ruan2024:quantumeyes; ruan2023:venus; wen2024:quantivine; ruan2023:vacsen; ruan2024:violet), we derive the following design principles to motivate our design decisions. Given the shared high-level objectives (QC interfaces), we use Ashktorab et al. (ashktorab2019:hqci)’s interview study to compare with the above challenges along with other QC education literature (wootton2021:interactiveTextbook; mykhailova2022:softwareQC; hu2024:qcLearning; tappert2019:experience; salehi2022:csQC). We also considered prior work regarding data science work pipelines (wang2024:supernova; crisan2021) given their similarity in terms of high-level procedures and intensive use of computational notebooks.
@@
(P1) Link conceptual ideas to quantum information. As specialized for processing quantum mechanical properties, QC programs exhibit more tight association of their inputs and outputs with low-level quantum information, compared to classical computing (ashktorab2019:hqci). Studies also imply that an understanding of rudimentary concepts like qubits and gates does not necessarily mean an ability to apply and interpret actual circuits based on those concepts (mykhailova2022:softwareQC; wootton2021:interactiveTextbook; ashktorab2019:hqci) (e.g., 64-qubit device requires $42^{6}$ 4 bits $=2,306$ petabytes). In our design iteration, for example, Jeng shared her experience in making custom classical programs to encode human-readable inputs like image files as quantum states for QML algorithms. Similarly, when learning Shor’s algorithm (shors:algorithm) that divides integers into two interger factors, Kim had struggles in understanding how the algorithm encodes integer multiplications as gates. This generalization challenge makes it difficult to conceptually express QC programs, which in turn complicates learning real-world applications like cryptography.
@@
On the other hand, QC program outputs are essentially the counts of measured qubit states that are expressed as bit strings (e.g., |$\ket{00101011}$). For instance, Kim needed to manually convert the output bit strings of Shor’s algorithm into human-readable numbers. Similarly, Jeng had to come up with custom functions to convert measurement outcomes (as probability distributions) into RGB images. Furthermore, a QC program with $n$ qubits can result in up to $2^{n}$ states, indicating the need for scalability in expressing QC output. For example, the previous Shor’s algorithm tutorial by Qiskit with a fixed factor of 15 requires measuring 8 qubits. This measurement results in $2562^{8}=$ 256 possible states on a real machine, where a histogram is not necessarily an ideal representation method due to the excessive visual density. Ashktorab et al. (ashktorab2019:hqci) also raise this scalability issue in quantum state and circuit visualizations because existing representation methods may fail to support interpreting circuits beyond a certain size. Therefore, usable QC interfaces must support mapping developers’ high-level ideas to low-level quantum information to improve flexibility in circuit composition and interpretability in result analysis.
@@
(P2) Provide support at different levels of computing. Through the design iterations, we recognized that QC developers would work on varying levels of computing, like architecture, compiler, algorithms, etc. For example, Smith wanted a lot of details in optimization outcomes while Kim did not. Jeng’s needs for optimization detail increased as she started working on architecture-related research. Similarly, Ashktorab et al. (ashktorab2019:hqci) identify different QC-related roles: scientists who want full access to low-level details; developers who do not necessarily want those in-depth information; and enthusiasts who are interested in exploring QC capabilities with easy interactions. These different needs for information reflect the varying scales and goals in the practice and research of QC.
@@
(P3) Apply usability standards to QC tools. During our design iterations, replication, sharing, and context switch were recurring, intertwined topics. Smith and her collaborators had to come up with custom methods to share circuits and outcomes in a reliable way, and Kim has experienced too basic documentations lacking examples and detailed descriptions. In this case, applying widely known usability techniques can empower QC researchers and practitioners. For instance, a common method for exporting and retrieving experiment setups for running circuits (e.g., benchmark codes) and the results (e.g., measurement outcomes) can better facilitate collaboration among QC researchers and practitioners. Given that understanding QC theory, scripting in Python, and comprehending dense code bases are different skills, QC tools need methods for navigating APIs, such as an overview dashboard. In doing so, a usable QC interface will need to minimize context switching by offering on-demand information at where developers are working because frequent context switching and the lack of rich documentation can negatively affect productivity in programming (myer2014; myers2016; shrestha2020).
@@ 6.1. Overall Interface Techniques
(T0a) Integration with computational notebooks. As noted earlier, tools and libraries for QC offer APIs to be used in computational notebooks. While prior work and existing tools mainly employ stand-alone interfaces, it is cumbersome to export and import data from those interfaces to computational notebooks (e.g., matching file paths, restructuring data). Thus, our prototypes integrate features for different tasks within the same working environment (P3: Baseline usability support), which we decided as computational notebooks given its wide adoption by current QC developers (ashktorab2019:hqci).
@@ 6.2. Techniques for Circuit Composition
(T1a) Problem-oriented circuit writing. When using most QC tools, developers have to specify a circuit using atomic gates like H, X, and CNOT gates except some subroutines that are mathematically well generalized, such as quantum Fourier transformation (QFT) (coppersmith2002:qft). While graphical circuit composers (ibmComposer; quirk) enables visual programming of QC circuits, they are limited to reduce this inherent gap between conceptual ideas and low-level circuit designs. Developers still have to specify low-level gates, and it is less scalable for circuits with more than a few qubits (ashktorab2019:hqci). When adopting templates for well-known algorithms, developers still need to update those templates, fix needy-greedy details, and deal with unexpected downstream effects. Instead, our prototype interface for circuit composition allows for choosing a problem to encode by incorporating prior methods individually proposed by QC researchers (P1: Link concepts and quantum), such as truth tables (revkit), image convolution (jeng2023:convolution). Once choosing a problem to solve, then developers can provide parameters using interpretable values like integers, image files, logical expressions, etc.
(T1b) Circuit verification. Writing a QC circuit involves technical constraints like setting the correct number of qubits to use and updating operation parameters. When editing circuits, for example, developers may fail to checking those constraints, which commonly happens in code recycling (mili1995:softwarereuse). In the best case, developers just need to debug them, whereas wrong results, the worst case, can cause a huge financial cost for redoing everything. Thus, our interface verifies a circuit while writing it, so that QC developers can flexibly write and revise their circuits with reduced concerns in technical constraints (P1: Link concept and quantum).
@@
(T1d) Easier access to QASM format. While QC developers use Open Quantum Assembly Language (or QASM) (qasm) as a common circuit representation method across different QC libraries, it is much harder to use than them. As basic usability support (P3), our interface allows for saving or copying QASM expressions for circuits written interactively.

### v3

@@ 3.1. Interfaces for Quantum Computing
Prior work in understanding how people use QC tools sets useful context for our work. By interviewing seven scientists in different domains, Ashktorab et al. (Ashktorab et al., 2019) report the needs for assisting beginners with using QC tools with limited mathematical knowledge. They also note that after passing a certain point in training, one might need and want to move on to more professional tools for a higher degree of freedom. Furthermore, they recognize the importance of sharing and replicating quantum programs as well as visualizing quantum information in a scalable way given that $n$ qubits result in $2^{n}$ states (worst case). Prior work based on education settings (Mykhailova and Svore, 2020; Hu et al., 2024; Tappert et al., 2019; Salehi et al., 2022) focused on the implementation of simpler circuits representing basic concepts, such as superposition and entanglement, and provided conceptual understanding of complex algorithms like quantum chemistry or optimization problems. In addition, by analyzing self-paced interactive textbook, Wootton et al. (Wootton et al., 2021) found that learners tended to stop at earlier chapters and frequently referred to appendix for linear algebra. These findings indicate two competing goals for QC support: introducing complex quantum algorithms in a conceptual way vs. training learners to work with low-level components of those algorithms. This work proposes and demonstrates approaches to conceptually representing components of quantum algorithms while not blocking the degree of freedom offered by low-level toolkits.
@@
Visualization takes an important role in making QC more understandable (Bethel et al., 2023), such as Bloch sphere (Bloch, 1946) for qubit states, circuit diagrams, and measurement histograms for QC output. On top of those common visual representations, prior approaches focused on interactive interfaces for circuit composition and result analysis. First, interactive circuit composers like IBM Quantum Composer (IBM, 2016) and Quirk (Gidney, 2016) let users drag and drop gates on a circuit visualization. Yet, those tools still require a deeper understanding of how each gate functions with respect to bigger conceptual ideas, which Classiq (Classiq, 2024) is trying to address with an interactive whiteboard (Figure 2A). In addition, to improve the interpretability of changes to qubits’ states along with operations in a circuit, prior work proposed approaches to integrating this information on a circuit diagram. For instance, QuFlow (Lin et al., 2018) shows possible states of qubits with their probabilities in number on a circuit, and QuantumEyes (Ruan et al., 2024a) shows them as an area chart with small diagrams explaining how each gate affects the probabilities of qubit states. VENUS (Ruan et al., 2023b) adopts geometrical probability representations for multi-qubit states given that traditional 3D representations like Bloch spheres are limited to show multi-qubit states effectively. However, showing qubit state probabilities on a circuit are feasible only when a circuit includes a few qubits because probability information is not tractable for circuits with more than a few qubits, which is why we need QC. Instead, Quantivine (Wen et al., 2024) abstracts a circuit representation in terms of structural patterns (e.g., repeated gates) and allows for filtering a few qubits for more details like their interactions with other qubits. Note that one can easily encode the structure of a circuit using Qiskit and check their decompositions. Taking one step further, our work proposes and implements interactive visualizations for comparing logical and physical circuits and inspecting how physical qubits’ status change over the execution of a circuit.
@@
In analyzing QC results, developers need to consider non-negligible amount of gate and readout errors as well as adapt the histogram-based outputs for specific problem types. To enable that, VACSEN (Ruan et al., 2023a) parallelly shows quantum circuits with the decoherence time and error rates of the gates and measurements involved in executing the circuits. In particular, VACSEN shows error information with timestamps because QC machines’ physical properties change over time. To better support QC-based machine learning (quantum machine learning or QML), VIOLET (Ruan et al., 2024b) offers a expert-targeted dashboard that shows QC outputs in the context of neural network (e.g., output states as features that a model has learned). Yet, prior work does not offer a direct way to represent QC outputs with error rates and/or in a human-understandable format, which our work aims to propose approaches to complementing.
@@
In adopting these complex computing approaches, developers frequently use computational notebooks (e.g., Jupyter Notebook) as they are easy to share, reaccess (e.g., compared to running them on a terminal), and refine (Wang et al., 2024). Computational notebooks allow for integrating custom widgets, making it easier to combine work pipelines (Wang et al., 2024; Harden et al., 2024). Yet, prior approaches in QC and complex computing methods tend to showcase standalone tools rather than those integrated with working environments like computational notebooks. In addition, they often lack support for tasks pertaining to scripting and sharing codes for using them. To address these integration-wise challenges, our work aims to propose interaction techniques for a wider range of QC lifecycle by integrating them within computational notebooks.
@@ 3.2. Interface Support for Data-intensive Work
Prior work has proposed interfaces for different tasks in data-intensive work pipelines, such as data analysis and statistical modeling. Given data work is often done on a Jupyter notebook environment, for example, prior work suggested widgets for various tasks like developing statistical models (Jun et al., 2022) and analyzing text data (Lam et al., 2024). Dashboard-based interfaces are also common for complex statistical analysis (Liu et al., 2021; Sarma et al., 2024) and literature analysis (Narechania et al., 2022). Those interfaces commonly offer descriptions regarding steps that a user is taking. On the other hand, current QC users need to use the same tool regardless of their tasks because many tasks are constrained by QPUs provided via those tools. For instance, gates defined using Cirq can be optimized only by Cirq because the optimization algorithms rely on QPUs accessed via Cirq. Furthermore, users need to frequently switch between the above described QC tasks, requiring interfaces overarching those tasks. Thus, our work also explores how to adapt prior user interface techniques across various QC tasks.
@@ 4.2. Identified Challenges
(C1) Difficulty in expressing ideas and interpreting results conceptually. A recurring topic during our design iterations was the unmatched connection between conceptual ideas and quantum information. In both theoretical and applied QC courses, for example, Kim learned Shor’s algorithm (Shor, 1994) for factoring a large natural number into two other numbers, which is expected to advance quantum cryptography. However, he learned only about theoretical proofs and potential application areas (which is common in many QC classes (Hu et al., 2024; Tappert et al., 2019)), but was not able to apply the algorithm to novel problems. It is tricky to try out different numbers for Shor’s algorithm because doing so requires more profound knowledge than what typical beginners would have. Qiskit provides a tutorial for how to do so [^5] https://github.com/qiskit-community/qiskit-community-tutorials/blob/master/algorithms/shor_algorithm.ipynb, but it only shows a case where one factor is fixed at 15. To apply a different factor, Kim had to come up with a method to encode that number using lots of X and Swap gates [^6] A Swap gate changes the states of two qubits, which was beyond Kim’s initial capability (c.f. (Beauregard, 2003)).
Similarly, when Jeng was developing image-convolution techniques for a QML project, she had to come up with a relatively large suite of custom codes to convert images to quantum states (and vice versa). Currently, QML developers lack shared standards that work generically with other image-processing algorithms. Instead, they often need to edit codes from tutorials, making it tedious to apply those codes for novel use cases. Even though more experienced in QC research, Smith was not necessarily able to immediately come up with those methods because Jeng focused on specific QML techniques while Smith was more interested scalable optimization techniques across different application scenarios.
@@ 5. Design Principles for QC Interfaces
Based on our iterative design process and prior work in QC interfaces (Ashktorab et al., 2019; Lin et al., 2018; Ruan et al., 2024a, 2023b; Wen et al., 2024; Ruan et al., 2023a, 2024b), we derive the following design principles to motivate our design decisions. Given the shared high-level objectives (QC interfaces), we use Ashktorab et al. (Ashktorab et al., 2019)’s interview study to compare with the above challenges along with other QC education literature (Wootton et al., 2021; Mykhailova and Svore, 2020; Hu et al., 2024; Tappert et al., 2019; Salehi et al., 2022). We also considered prior work regarding data science work pipelines (Wang et al., 2024; Crisan et al., 2021) given their similarity in terms of high-level procedures and intensive use of computational notebooks.
@@
(P1) Link conceptual ideas to quantum information. As specialized for processing quantum mechanical properties, QC programs exhibit more tight association of their inputs and outputs with low-level quantum information, compared to classical computing (Ashktorab et al., 2019). Studies also imply that an understanding of rudimentary concepts like qubits and gates does not necessarily mean an ability to apply and interpret actual circuits based on those concepts (Mykhailova and Svore, 2020; Wootton et al., 2021; Ashktorab et al., 2019). In our design iteration, for example, Jeng shared her experience in making custom classical programs to encode human-readable inputs like image files as quantum states for QML algorithms. Similarly, when learning Shor’s algorithm (Shor, 1994) that divides integers into two interger factors, Kim had struggles in understanding how the algorithm encodes integer multiplications as gates. This generalization challenge makes it difficult to conceptually express QC programs, which in turn complicates learning real-world applications like cryptography.
@@
On the other hand, QC program outputs are essentially the counts of measured qubit states that are expressed as bit strings (e.g., $|{00101011}\rangle$). For instance, Kim needed to manually convert the output bit strings of Shor’s algorithm into human-readable numbers. Similarly, Jeng had to come up with custom functions to convert measurement outcomes (as probability distributions) into RGB images. Furthermore, a QC program with $n$ qubits can result in up to $2^{n}$ states, indicating the need for scalability in expressing QC output. For example, the previous Shor’s algorithm tutorial by Qiskit with a fixed factor of 15 requires measuring 8 qubits. This measurement results in $2562^{8}=$ 256 possible states on a real machine, where a histogram is not necessarily an ideal representation method due to the excessive visual density. Ashktorab et al. (Ashktorab et al., 2019) also raise this scalability issue in quantum state and circuit visualizations because existing representation methods may fail to support interpreting circuits beyond a certain size. Therefore, usable QC interfaces must support mapping developers’ high-level ideas to low-level quantum information to improve flexibility in circuit composition and interpretability in result analysis.
@@
(P2) Provide support at different levels of computing. Through the design iterations, we recognized that QC developers would work on varying levels of computing, like architecture, compiler, algorithms, etc. For example, Smith wanted a lot of details in optimization outcomes while Kim did not. Jeng’s needs for optimization detail increased as she started working on architecture-related research. Similarly, Ashktorab et al. (Ashktorab et al., 2019) identify different QC-related roles: scientists who want full access to low-level details; developers who do not necessarily want those in-depth information; and enthusiasts who are interested in exploring QC capabilities with easy interactions. These different needs for information reflect the varying scales and goals in the practice and research of QC.
@@
(P3) Apply usability standards to QC tools. During our design iterations, replication, sharing, and context switch were recurring, intertwined topics. Smith and her collaborators had to come up with custom methods to share circuits and outcomes in a reliable way, and Kim has experienced too basic documentations lacking examples and detailed descriptions. In this case, applying widely known usability techniques can empower QC researchers and practitioners. For instance, a common method for exporting and retrieving experiment setups for running circuits (e.g., benchmark codes) and the results (e.g., measurement outcomes) can better facilitate collaboration among QC researchers and practitioners. Given that understanding QC theory, scripting in Python, and comprehending dense code bases are different skills, QC tools should enable navigating APIs for different tasks, for example, on an overview dashboard. In doing so, a usable QC interface will need to minimize context switching by offering on-demand information at where developers are working because frequent context switching and the lack of rich documentation can negatively affect productivity in programming (Meyer et al., 2014; Myers et al., 2016a; Shrestha et al., 2020).
@@ 6.1. Overall Interface Techniques
(T0a) Integration with computational notebooks. As noted earlier, tools and libraries for QC offer APIs to be used in computational notebooks. While prior work and existing tools mainly employ stand-alone interfaces, it is cumbersome to export and import data from those interfaces to computational notebooks (e.g., matching file paths, restructuring data). Thus, our prototypes integrate features for different tasks within the same working environment (P3: Baseline usability support), which we decided as computational notebooks given its wide adoption by current QC developers (Ashktorab et al., 2019).
@@ 6.2. Techniques for Circuit Composition
(T1a) Problem-oriented circuit writing. When using most QC tools, developers have to specify a circuit using atomic gates like H, X, and CNOT gates except some subroutines that are mathematically well generalized, such as quantum Fourier transformation (QFT) (Coppersmith, 2002). While graphical circuit composers (IBM, 2016; Gidney, 2016) enables visual programming of QC circuits, they are limited to reduce this inherent gap between conceptual ideas and low-level circuit designs. Developers still have to specify low-level gates, and it is less scalable for circuits with more than a few qubits (Ashktorab et al., 2019). When adopting templates for well-known algorithms, developers still need to update those templates, fix needy-greedy details, and deal with unexpected downstream effects. Instead, our prototype interface for circuit composition allows for choosing a problem to encode by incorporating prior methods individually proposed by QC researchers (P1: Link concepts and quantum), such as truth tables (Soeken, 2019), image convolution (Jeng et al., 2023). Once choosing a problem to solve, then developers can provide parameters using interpretable values like integers, image files, logical expressions, etc.
(T1b) Circuit verification. Writing a QC circuit involves technical constraints like setting the correct number of qubits to use and updating operation parameters. When editing circuits, for example, developers may fail to checking those constraints, which commonly happens in code recycling (Mili et al., 1995). In the best case, developers just need to debug them, whereas wrong results, the worst case, can cause a huge financial cost for redoing everything. Thus, our interface verifies a circuit while writing it, so that QC developers can flexibly write and revise their circuits with reduced concerns in technical constraints (P1: Link concept and quantum).
@@
(T1d) Easier access to QASM format. While QC developers use Open Quantum Assembly Language (or QASM) (Cross et al., 2017) as a common circuit representation method across different QC libraries, it is difficult to use them. As basic usability support (P3), our interface allows for saving or copying QASM expressions for circuits written interactively along with Python codes.